The supplement of an angle measures four times more than measure of its complement. Determine the measure of the complement of this angle.
Answer by Alan3354(21581)

The supplement of an angle measures four times more than measure of its complement. Determine the measure of the complement of this angle.
------
180-x = 4(90-x) = 360-4x
3x = 180
x = 60º
Complement = 90-60
= 30º